# Curvature and local matchings of conference graphs and extensions

### v2

@@ Abstract
We confirm a conjecture of Bonini et al. on the precise Lin–Lu–Yau curvature values of conference graphs, i.e., strongly regular graphs with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, with $\gamma\geq 2$. Our method only depends on the parameter relations and applies to more general classes of amply regular graphs. In particular, we develop a new combinatorial method for showing the existence of local perfect matchings. A key observation is that counting common neighbors leads to useful quadratic polynomials. Our result also leads to an interesting number theoretic consequence on quadratic residues.
@@ 1 Introduction
The study of discrete notions of Ricci curvature has attracted much attention and found important applications in graph theory, see, e.g., [2, 3, 10, 21, 23, 24, 28, 29]. Ricci curvature is a key concept in Riemannian geometry and geometric analysis. The introduction of analogous concepts of Ricci curvature on graphs brings new deep insights and tools to the study of graph theory. In [28, 29], Ollivier introduced the concept of coarse Ricci curvature for Markov chains on metric spaces, including graphs. In [23], Lin, Lu, and Yau introduced a modified version of Ollivier’s coarse Ricci curvature on graphs. This concept of modified curvature, which we call the Lin–Lu–Yau curvature, has been proved to be very useful to detect geometric and analytic properties of the underlying graphs [14, 13, 12, 25, 26, 27].
Explicit calculations of Lin–Lu–Yau curvature on graphs with local regularity restrictions have important applications in the study of locally correctable and some locally testable binary linear codes [20], and in bounding the diameter and eigenvalues of amply regular graphs [22, 19]. A $d$-regular graph $G$ with $n$ vertices is called an amply regular graph with parameters $(n,d,\alpha,\beta)$ if any two adjacent vertices have $\alpha$ common neighbors and any two vertices of distance $2$ have $\beta$ common neighbors. Amply regular graphs with diameter $2$ are strongly regular. For any two adjacent vertices, denote by $\kappa(x,y)$ the Lin–Lu–Yau curvature of the edge $xy$. In the works of [22, 19], it has been shown that for any two adjacent vertices $x,y$ of an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$, the following lower bound estimates hold:
  • If $\beta=\alpha>1$, then $\kappa(x,y)\geq\frac{2}{d}$.
  • If $\beta>\alpha\geq 1$, then $\kappa(x,y)\geq\frac{3}{d}$.
The Lin–Lu–Yau curvature of amply regular graphs with $\alpha>\beta\geq 2$ is rather mysterious. It is still open whether such graphs have positively lower bounded Lin–Lu–Yau curvature or not.
@@
Let $G=(V,E)$ be a $d$-regular graph. For any edge $xy\in E$, we have
@@
The crucial role of local matching conditions in the curvature calculation has been first observed in [3, 32], and further developed in [5, 15, 27]. For an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$, the above upper bound reads $\kappa(x,y)\leq\frac{2+\alpha}{d}$. A natural question is for which parameters this upper bound is achieved.
Bonini et al. [5] derived formulas for the Lin–Lu–Yau curvature of strongly regular graphs based on the parameters of graphs and the size of a maximal matching in the core neighborhood. They further proposed the following conjecture [5, Conjecture 1.7].
@@
Let $G=(V,E)$ be a strongly regular conference graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, with $\gamma\geq 2$. Then the Lin–Lu–Yau curvature satisfies
@@
Recall that a conference graph is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$. Conference graphs have very interesting properties: While being highly symmetric, they are quasirandom [1, Theorem 9.3.1]. The existence of conference graphs are known for small values of $\gamma$ and for each $\gamma$ such that $14\gamma+$ 1 is a prime power (Paley graphs). Let $n=4\gamma+1$ be a prime power and $GF(n)$ be the unique finite field of order $n$. Recall that a Paley graph $P(n)$ of order $n$ is a graph on $n$ vertices with two vertices adjacent if their difference is a non-zero square in $GF(n)$. However, there are still many values of $\gamma$ for which the existence of conference graphs is open. Observe by Theorem 1.1 that the Lin–Lu–Yau curvature of a conference graph is no greater than $\frac{2+(\gamma-1)}{2\gamma}=\frac{1}{2}+\frac{1}{2\gamma}$. Therefore, Conjecture 1.2 claims the upper bound is achieved for conference graphs.
We stress that, in general, the Lin–Lu–Yau curvature of an amply regular graph is not necessarily uniquely determined by its parameters $(n,d,\alpha,\beta)$. For instance, the $4\times 4$ Rook’s graph and the Shrikhande graph are both strongly regular with parameters $(16,6,2,2)$, but Bonini et al. [5] calculated their Lin–Lu–Yau curvatures to be $\kappa=\frac{2}{3}$ and $\kappa=\frac{1}{3}$, respectively. In particular, the $4\times 4$ Rook’s graph takes the upper bound of the Lin–Lu–Yau curvature, while the Shrikhande graph does not.
Huang, Liu, and Xia showed [19, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ for any amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $d\leq 2\beta-\alpha-1$. However, the parameters of conference graphs violate the restriction $d\leq 2\beta-\alpha-1$. The lower curvature bound of amply regular graph with $\beta>\alpha\geq 1$ in [19], which has been mentioned earlier, applies to conference graphs and provides a positive lower bound $\frac{3}{d}$.
After finishing the first version of this paper, we notice from a very recent preprint [6] that the Lin–Lu–Yau curvature of the special case of Paley graphs $P(n)$ of even power order $n=p^{2m}$ has been confirmed to be $\frac{2+\alpha}{d}$ in an unpublished manuscript by Bonini, Liu, Semmens and Tran. Very interesting extensions to generalized Paley graphs have been carried out in [6].
Our main result in this paper confirms the Conjecture 1.2.
@@
Indeed, our method depends only on the parameter relations and applies to more general classes of amply regular graphs. Below, we provide several sufficient parameter conditions.
@@
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Suppose that one of the following conditions holds:
  • (1)
    $\beta=\alpha+1$, $n<3d-2\alpha$ and $d>\alpha+\sqrt{6\alpha+\frac{1}{4}}+\frac{3}{2}$;
  • (2)
    $\beta=\alpha+2$, $n<3d-2\alpha$ and $d\geq\alpha+\sqrt{10\alpha+\frac{41}{4}}+\frac{3}{2}$;
  • (3)
    $\beta>\alpha+2$, $n<3d-2\alpha$ and $\left|d-\frac{3}{2}\beta\right|\geq\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20}$;
  • (4)
    $\beta\geq\frac{2\sqrt{3}}{3}\alpha+7$ and $n<3d-2\alpha$;
  • (5)
[…]
Then the Lin–Lu–Yau curvature satisfies
$\kappa(x,y)=\frac{2+\alpha}{d},\,\,\text{for all}\,\,xy\in E.$
By Theorem 1.1, Theorem 1.4 is equivalent to the following result in terms of local matchings.
Theorem 1.4.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Suppose that one of the $5$ parameter conditions in Theorem 1.4 holds. Then there exists a perfect matching between $N_{x}$ and $N_{y}$ for all $xy\in E$.
Interestingly, the parameter conditions stated in Theorem 1.4 are fulfilled by many known strongly regular graphs, such as Paley graphs, McLaughlin graph and so on. We provide a list of more graphs (on at most $512$ vertices) satisfying one of the parameter conditions in Theorem 1.4 in the Appendix. In particular, by the condition $(1)$ of Theorem 1.4, Conjecture 1.2 holds true for $\gamma>6$. The condition $(5)$ applies to many graphs of the mysterious case of amply regular graphs with $\alpha>\beta$, see the Appendix.
Our proof provides a new combinatorial method for showing the existence of local perfect matchings. This new trick can also be employed to deal with many cases of amply regular graphs other than those in Theorem 1.4. For example, the restriction $n<3d-2\alpha$ is not always necessary. We refer to Remark 3.4 and Example 3.5 for more detailed discussions.
Applying Theorem 1.3 to Paley graphs, we obtain an interesting result of number theory. Theorem 1.3 tells that there exists a perfect matching between $N_{x}$ and $N_{y}$ for any adjacent vertices $x,y$ of a Paley graph $P(n)$ with $n>5$. This leads to the following consequence.
@@
Let $n>5$ be a prime power congruent to $1$ modulo $4$. Consider two numbers $x,y\in GF(n)$ such that $x-y$ is a non-zero square in $GF(n)$. For any subset $S\subset GF(n)\setminus\{x,y\}$ with $|S|\geq\frac{3}{4}(n-1)$, there exist $w,z\in S$ such that $x-w,w-z,z-y$ are non-zero squares and $x-z,y-w$ are not squares in $GF(n)$.
In particular, this corollary applies to quadratic residues when $n>5$ is a Pythagorean prime, i.e., a prime congruent to $1$ modulo $4$.
Furthermore, we study another upper bound of the Lin–Lu–Yau curvature, which is given in terms of the Laplacaian eigenvalue and is an analogue to the Lichnerowicz Theorem in Riemannian geometry.
Theorem 1.6 (Discrete Lichnerowicz Theorem [28, 23]).
Let $G=(V,E)$ be a finite connected graph. Then the Lin–Lu–Yau curvature statisfies
$\min_{xy\in E}\kappa(x,y)\leq\lambda_{2},$
where $\lambda_{2}$ is the first nonzero eigenvalue of the normalized Laplacian $L$. Here the normalized Laplacian $L$ is defined as $L=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, with $A$ being the adjacency matrix of the graph $G$ and $D$ being the diagonal matrix of vertex degrees.
Following [14], we say a graph $G$ is Lichnerowicz sharp if $\lambda_{2}=\min_{xy\in E}\kappa(x,y)$, i.e., the above upper bound is achieved. We observe the following interesting phenomena.
Theorem 1.7.
Let $G=(V,E)$ be a connected strongly regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$. If $G$ is Lichnerowicz sharp, i.e., $\min_{xy\in E}\kappa(x,y)=\lambda_{2}$, then
$\kappa(x,y)=\frac{2+\alpha}{d}.$
Indeed, all distance regular graphs with the two curvature upper bounds in Theorem 1.1 and Theorem 1.6 both achieved, i.e.,
$\min_{xy\in E}\kappa(x,y)=\lambda_{2}=\frac{2+\alpha}{d},$
have been fully classified in [14, Theorem 6.4]. Accordingly, we obtain the following classification of Licherowicz sharp strongly regular graphs with $\beta\geq\alpha$.
Corollary 1.8.
The connected Lichnerowicz sharp strongly regular graphs with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$ are precisely the following ones: The cocktail party graphs $CP(k),k\geq 2$, the Paley graph $P(9)$, the Johnson graphs $J(5,2)$ and $J(6,2)$ and the demi-cube $Q^{5}_{(2)}$ (also called Clebsch graph).
The rest of the our paper is organized as follows. In Section 2, we recall definitions and properties about the Lin–Lu–Yau curvature, perfect matching, and strongly/amply regular graphs. In Section 3, we prove Theorem 1.3, Theorem 1.4 and Corollary 1.5, and present possible ways to generalize our conclusion. In Section 4, we prove the Theorem 1.7 and Corollary 1.8.
@@ 2.1 Lin–Lu–Yau curvature and local matchings
$\mu_{x}^{p}(y)=\left\{\begin{array}[]{ll}p,&\hbox{if $y=x$;}\\ \frac{1-p}{\mathrm{deg}(x)},&\hbox{if $y\sim x$;}\\ 0,&\hbox{otherwise,}\end{array}\right.$
where $\mathrm{deg}(x):=\sum_{y\in V:y\sim x}1$ is the vertex degree of $x$.
@@
Notice that $\kappa_{1}(x,y)$ is always 00. Hence, the Lin–Lu–Yau curvature $\kappa(x,y)$ equals the negative of left derivative of the function $p\mapsto\kappa_{p}(x,y)$ at $p=1$.
Boure et al. [7] proved the following relation between $p$-Ollivier curvature and Lin–Lu–Yau curvature for an edge $xy\in E$ of a $d$-regular graph $G=(V,E)$:
@@
This relation leads in particular to the following expression of Lin–Lu–Yau curvature of a $d$-regular graph:
@@
Hall’s marriage Theorem is an fundamental tool for proving Conjecture 1.2.
@@
holds, where $\Gamma_{V_{2}}(S):=\{v\in V_{2}|\,\,\text{there exists }\,w\in S\,\text{such that}\,\,vw\in E\}$.
@@ 2.2 Strongly/Amply regular graph
Let $G$ be a $d$-regular graph with $n$ vertices which is neither complete nor empty. Then $G$ is called call a strongly regular graph with parameters $(n,d,\alpha,\beta)$ if:
@@
[…]
@@
    Any two vertices with distance $2$ have $\beta$ common neighbors.
The following theorem shows a relationship between the parameters of an amply regular graph.
@@
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. then
@@
Let $M$ be the number of pairs of ($(\{v_{1},v_{2}\},u)$, where $v_{1},v_{2},u\in V$ and $u$ is a common neighbour of $v_{1}$ and $v_{2}$. For any $u\in V$, there are $\binom{d}{2}$ pairs of such $\{v_{1},v_{2}\}$. Thus, $M=\binom{d}{2}n$. For any two adjacent vertices $v_{1}$ and $v_{2}$, there are $\alpha$ such $u$. For any two non-adjacent vertices $v_{1}$ and $v_{2}$, there are at either $\beta$ or 00 such $u$. Therefore,
@@
where $e=|E|=\frac{nd}{2}$, and the inequality (2.3) follows. It’s clear that the equality holds if and only if $G$ is strongly regular. ∎
3 Proof of Theorem 1.3, Theorem 1.4 and Corollary 1.5
In this section, we prove Theorem 1.3, Theorem 1.4 and Corollary 1.5. First, we observe the following lemma which is useful in applying Hall’s marriage Theorem (Theorem 2.4).
@@
Let $G=(V,E)$ be an amply regular graphs with parameters $(n,d,\alpha,\beta)$ such that $d\geq\alpha+2$. We have the following properties:
  • (i)
    If $n<3d-2\alpha$, then $d\leq 2\beta$.
  • (ii)
    If $2n<5d-3\alpha$, then $d\leq\frac{3}{2}\beta$.
@@
Since $d\geq\alpha+2$, we deduce $d\leq 2\beta$. This proves $(i)$. $(ii)$ is deduced similarly. ∎
@@
The subsets above are defined by $\Delta_{xy}:=\Gamma(x)\cap\Gamma(y)$, $N_{x}:=\Gamma(x)\backslash(\Delta_{xy}\cup\{y\})$, $N_{y}:=\Gamma(y)\backslash(\Delta_{xy}\cup\{x\})$, and $P_{xy}:=V\backslash(\Gamma(x)\cup\Gamma(y))$, where $\Gamma(x):=\{z\in V|\,xz\in E\}$, and $\Gamma(y):=\{z\in V|\,yz\in E\}$. It is direct to check by definition that
@@
Since $y$ and $v$ have $\beta$ common neighbors, $v$ has $\beta-1-\ell$ neighbors in $\Delta_{xy}$. Since $x$ and $v$ have $\alpha$ common neighbors, $v$ has $\alpha-\beta+1+\ell$ neighbors in $N_{x}$. Since $v$ has $d-1$ neighbors except $x$, the number of neighbors of $v$ in $P_{xy}$ is $d-\alpha-1-\ell$. ∎
Now, we are prepared to prove Theorem 1.4.
@@
Let $G=(V,E)$ be an amply regular graph satisfying one of the conditions (1)–(5), and $x,y\in V$ be two adjacent vertices. Recall from [19, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $d\leq 2\beta-\alpha-1$. In the following proof, we assume that $d\geq 2\beta-\alpha$.
We claim that each of the conditions (1)–(5) yields $d>\alpha+2$. Indeed, it is clear that conditions (1) and (2) yield $d>\alpha+2$. Under the condition (3) or (4), $d\geq\beta>\alpha+2$. Under the condition (5), $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}>\alpha+3$. Therefore, the claim holds. Recalling (3.2), we derive $|N_{x}|=|N_{y}|=d-\alpha-1>1$. That is, both $N_{x}$ and $N_{y}$ contain at least two vertices.
Recall from [22, Theorem 1.5] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $\alpha=0$ and $\beta\geq 2$. If $\alpha=0$, then each of the conditions (2)–(4) leads to $\beta\geq 2$ and hence $\kappa(x,y)=\frac{2}{d}$. Under the condition (1), $\alpha=0$ implies $\beta=1$ and $d>2$. However, by the condition $n<3d-2\alpha$ and Lemma 3.2, we have $d\leq 2\beta=2$, which is a contradiction. That is, under the condition (1), the case $\alpha=0$ is not allowed. Under the condition (5), it’s clear that the case $\alpha=0$ can not happen. In the following proof, we assume that $\alpha>0$.
@@
always holds. For each vertex $v_{i}\in S:=\{v_{1},v_{2},\ldots,v_{b}\}$, let $x_{i}$ be the number of neighbors of $v_{i}$ in $N_{y}$. By Lemma 3.3, the vertex $v_{i}$ has $d-\alpha-1-x_{i}$ neighbors in $P_{xy}$. If $|P_{xy}|=0$, then we have $x_{i}=d-\alpha-1=|N_{y}|$. Therefore, the inequality (3.3) clearly holds. In the following proof, we assume that $|P_{xy}|>0$, i.e., $n-2d+\alpha>0$.
@@
We argue by contradiction. Suppose $|T|\leq b-1$.

### v3

@@ Abstract
We prove a conjecture of Bonini et al. on the precise values of the Lin–Lu–Yau curvature of conference graphs, i.e., strongly regular graphs with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$. Our method depends only on the parameter relations and applies to broader classes of amply regular graphs. In particular, we develop a new combinatorial approach to show the existence of local perfect matchings. A key observation is that counting common neighbors leads to useful quadratic polynomials. As a corollary, we derive an interesting number-theoretic result concerning quadratic residues.
@@ 1 Introduction
The study of discrete notions of Ricci curvature has attracted much attention and found important applications in graph theory, see, e.g., [2, 3, 10, 19, 21, 22, 26, 27]. Ricci curvature is a key concept in Riemannian geometry and geometric analysis. The introduction of analogous concepts of Ricci curvature on graphs brings new deep insights and tools to the study of graph theory. In [26, 27], Ollivier introduced the concept of coarse Ricci curvature for Markov chains on metric spaces, including graphs. In [21], Lin, Lu, and Yau introduced a modified version of Ollivier’s coarse Ricci curvature on graphs. This modified curvature, which we call the Lin–Lu–Yau curvature, has proved very useful in detecting geometric and analytic properties of the underlying graphs [12, 13, 14, 23, 24, 25, 28].
Explicit calculations of Lin–Lu–Yau curvature on graphs with local regularity restrictions have important applications in the study of locally correctable and some locally testable binary linear codes [18], and in bounding the diameter and eigenvalues of amply regular graphs [17]. A $d$-regular graph $G$ with $n$ vertices is called an amply regular graph with parameters $(n,d,\alpha,\beta)$ if any two adjacent vertices have $\alpha$ common neighbors and any two vertices at distance $2$ have $\beta$ common neighbors. Amply regular graphs with diameter $2$ are called strongly regular graphs. Denote by $\kappa(x,y)$ the Lin–Lu–Yau curvature of an edge $xy$. It is shown [17] that, for any edge $xy$ of an amply regular graph $G$ with parameters $(n,d,\alpha,\beta)$ with $\beta>\alpha\geq 1$, we have $\kappa(x,y)\geq\frac{3}{d}$.
@@
[14] Let $G=(V,E)$ be a $d$-regular graph. For any edge $xy\in E$, we have
@@
The crucial role of local matching conditions in the curvature calculation was first observed in [3, 29], and further developed in [5, 15, 25]. For an amply regular graph with parameters $(n,d,\alpha,\beta)$, the above upper bound reads $\kappa(x,y)\leq\frac{2+\alpha}{d}$. A natural question is for which parameters this upper bound is achieved.
Bonini et al. [5] derived formulas for the Lin–Lu–Yau curvature of strongly regular graphs in terms of the graph parameters and the size of a maximal matching in the core neighborhood. They further proposed the following conjecture [5, Conjecture 1.7].
@@
[5] Let $G=(V,E)$ be a strongly regular conference graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$ with $\gamma\geq 2$. Then the Lin–Lu–Yau curvature satisfies
@@
Recall that a conference graph is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$. Conference graphs have very interesting properties: While being highly symmetric, they are quasirandom [1]. The existence of conference graphs is known for each $\gamma$ such that $4\gamma+1$ is a prime power (i.e. Paley graphs). Let $n=4\gamma+1$ be a prime power and $GF(n)$ be the unique finite field of order $n$. Recall that a Paley graph $P(n)$ of order $n$ is a graph on $n$ vertices with two vertices adjacent iff their difference is a non-zero square in $GF(n)$. By Theorem 1.1, the Lin–Lu–Yau curvature of a conference graph is no greater than $\frac{2+(\gamma-1)}{2\gamma}=\frac{1}{2}+\frac{1}{2\gamma}$. Therefore, Conjecture 1.2 in fact claims that the upper bound is achieved for conference graphs.
We stress that, in general, the Lin–Lu–Yau curvature of an amply regular graph is not necessarily uniquely determined by its parameters. For instance, the $4\times 4$ Rook’s graph and the Shrikhande graph are both strongly regular with parameters $(16,6,2,2)$. However, Bonini et al. [5] computed their Lin–Lu–Yau curvatures to be $\frac{2}{3}$ and $\frac{1}{3}$, respectively. In particular, the $4\times 4$ Rook’s graph attains the upper bound of the Lin–Lu–Yau curvature, while the Shrikhande graph does not.
Huang, Liu, and Xia showed [17, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ for any amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $d\leq 2\beta-\alpha-1$. However, the parameters of conference graphs, namely $(4\gamma+1,2\gamma,\gamma-1,\gamma)$, do not satisfy this condition.
A recent work by Bonini, Liu, Semmens, and Tran [6] proves Conjecture 1.2 for Paley graphs $P(n)$ of even power order $n=p^{2m}$. Very interesting extensions to generalized Paley graphs have also been carried out in [6].
In this paper, we provide several sufficient conditions for the curvature of an amply regular graph to reach the natural upper bound. As a corollary, we prove Conjecture 1.2.
Theorem 1.3.
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$ such that $n<3d-2\alpha$. Suppose that one of the following conditions holds:
  • (1)
    $\beta=\alpha+1$ and $d>\alpha+\sqrt{6\alpha+\frac{1}{4}}+\frac{3}{2}$;
  • (2)
    $\beta\geq\alpha+2$ and $d\geq\frac{3}{2}\beta+\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+28\beta}$;
  • (3)
    $\beta\geq\frac{2\sqrt{3}}{3}\alpha+7$.
Then the Lin–Lu–Yau curvature satisfies
$\kappa(x,y)=\frac{2+\alpha}{d},\,\,\text{for all}\,\,xy\in E.$
In particular, condition (1) of Theorem 1.3 covers all conference graphs with $\gamma>6$. The cases for $2\leq\gamma\leq 6$ are verified separately in Section 3.
@@
By Theorem 1.1, proving Theorem 1.3 reduces to establishing the following result in terms of local matchings, which is equivalent to Theorem 1.3.
@@
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. If one of the $3$ parameter conditions in Theorem 1.3 holds, then there exists a perfect matching between $N_{x}$ and $N_{y}$ for every $xy\in E$.
Besides conference graphs, our parameter conditions are also fulfilled by many other strongly regular graphs, as shown in the arXiv version of this paper (see [9, Appendix]). In fact, we provide five more refined conditions in [9, Theorem 1.4], which generalize the three conditions in Theorem 1.3. For brevity, we do not include them here.
Our proof provides a new combinatorial method for showing the existence of local perfect matchings, which is also applicable to many cases of amply regular graphs beyond those in Theorem 1.3 and [9, Theorem 1.4]. For example, the restriction $n<3d-2\alpha$ is not always necessary. We refer to Remark 3.4 and Example 3.5 for more detailed discussions.
Applying Theorem 1.4 to Paley graphs yields an interesting number-theoretic result. Theorem 1.4 tells that there exists a perfect matching between $N_{x}$ and $N_{y}$ for any adjacent vertices $x,y$ of a Paley graph $P(n)$ with $n>5$. This leads to the following corollary.
@@
Let $n>5$ be a prime power congruent to $1$ modulo $4$. Consider two elements $x,y\in GF(n)$ such that xyx-y is a non-zero square in $GF(n)$. For any subset $S\subset GF(n)\setminus\{x,y\}$ with $|S|\geq\frac{3}{4}(n-1)$, there exist $w,z\in S$ such that xw,wz,zyx-w,w-z,z-y are non-zero squares and xz,ywx-z,y-w are not squares in $GF(n)$.
In particular, the above corollary applies to quadratic residues when $n$ is a Pythagorean prime (i.e., a prime congruent to $1$ modulo $4$) greater than $5$.
The rest of this paper is organized as follows. In Section 2, we recall definitions and properties about the Lin–Lu–Yau curvature, perfect matching, and strongly/amply regular graphs. In Section 3, we prove Theorem 1.3, Theorem 1.4, and Corollary 1.5, and present possible ways to generalize our conclusion.
@@ 2.1 Lin–Lu–Yau curvature and local matchings
$\mu_{x}^{p}(y)=\left\{\begin{array}[]{ll}p,&\hbox{if $y=x$;}\\ \frac{1-p}{\mathrm{deg}(x)},&\hbox{if $yx\in E$;}\\ 0,&\hbox{otherwise,}\end{array}\right$.
where $\mathrm{deg}(x)$ is the degree of the vertex $x$.
@@
Notice that $\kappa_{1}(x,y)$ is always 0. Hence, the Lin–Lu–Yau curvature $\kappa(x,y)$ is equal to the negative of the left derivative of the function $p\mapsto\kappa_{p}(x,y)$ at $p=1$.
Bourne et al. [7] proved the following relation between $p$-Ollivier curvature and Lin–Lu–Yau curvature for an edge $xy$ of a $d$-regular graph $G$:
@@
By the property of the Monge problem, this relation leads in particular to the following expression of Lin–Lu–Yau curvature of a $d$-regular graph:
@@
Hall’s marriage Theorem is a fundamental tool for proving Conjecture 1.2.
@@
where $\Gamma_{V_{2}}(S):=\{v\in V_{2}|\,\,\text{there exists }\,w\in S\,\text{such that}\,\,vw\in E\}$.
@@ 2.2 Strongly/Amply regular graph
Let $G$ be a $d$-regular graph with $n$ vertices which is neither complete nor empty.
The graph $G$ is called a strongly regular graph with parameters $(n,d,\alpha,\beta)$ if:
@@
The graph $G$ is called an amply regular graph with parameters $(n,d,$ $\alpha,\beta)$ if:
@@
    Any two vertices at distance two have $\beta$ common neighbors.
The following theorem shows a basic relationship between the parameters of an amply regular graph.
@@
Let $G=(V,E)$ be an amply regular graph with parameters $(n,d,\alpha,\beta)$. Then,
@@
Let $M$ be the number of pairs of $(\{v_{1},v_{2}\},u)$, where $v_{1},v_{2},u\in V$ and $u$ is a common neighbour of $v_{1}$ and $v_{2}$. For any $u\in V$, there are $\binom{d}{2}$ pairs of such $\{v_{1},v_{2}\}$. Thus, $M=\binom{d}{2}n$. For any two adjacent vertices $v_{1}$ and $v_{2}$, there are $\alpha$ such $u$. For any two non-adjacent vertices $v_{1}$ and $v_{2}$, there are at either $\beta$ or 0 such $u$. Therefore,
@@
where $e=|E|=\frac{nd}{2}$, and the inequality (2.3) follows. It is clear that the equality holds if and only if $G$ is strongly regular. ∎
3 Proof of Theorem 1.3, Theorem 1.4, and Corollary 1.5
In this section, we prove Theorem 1.3, Theorem 1.4, and Corollary 1.5. First, we observe the following lemma which is useful in applying Hall’s marriage Theorem (Theorem 2.4).
@@
Let $G=(V,E)$ be an amply regular graphs with parameters $(n,d,\alpha,\beta)$ such that $d\geq\alpha+2$. If $n<3d-2\alpha$, then $d\leq 2\beta$.
@@
Since $d\geq\alpha+2$, we deduce $d\leq 2\beta$. ∎
@@
The subsets above are defined by $\Delta_{xy}:=\Gamma(x)\cap\Gamma(y)$, $N_{x}:=\Gamma(x)\backslash(\Delta_{xy}\cup\{y\})$, $N_{y}:=\Gamma(y)\backslash(\Delta_{xy}\cup\{x\})$, and $P_{xy}:=V\backslash(\Gamma(x)\cup\Gamma(y))$, where $\Gamma(x):=\{z\in V|\,xz\in E\}$ and $\Gamma(y):=\{z\in V|\,yz\in E\}$. It is direct to check by definition that
@@
Since $y$ and $v$ have $\beta$ common neighbors, $v$ has $\beta-1$-$\ell$ neighbors in $\Delta_{xy}$. Since $x$ and $v$ have $\alpha$ common neighbors, $v$ has $\alpha-\beta+1+\ell$ neighbors in $N_{x}$. Since $v$ has d1d-1 neighbors except $x$, the number of neighbors of $v$ in $P_{xy}$ is $d-\alpha-1$-$\ell$. ∎
We now proceed to the proof of Theorem 1.3.
@@
Let $G=(V,E)$ be an amply regular graph satisfying one of the conditions (1)–(3), and let $xy\in E$ be an edge. Recall from [17, Proposition 3.6] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $d\leq 2\beta-\alpha-1$. Thus, we assume throughout the proof that $d\geq 2\beta-\alpha$.
We claim that each of the conditions (1)–(3) yields $d>\alpha+2$. Indeed, it is clear that condition (1) yields $d>\alpha+2$. Under the condition (2), $d\geq\frac{3}{2}\beta>\alpha+2$. Since $d\geq\beta$, condition (3) implies $d>\alpha+2$. Therefore, the claim holds. Recalling (3.2), we derive $|N_{x}|=|N_{y}|=d-\alpha-1>1$. That is, both $N_{x}$ and $N_{y}$ contain at least two vertices.
Recall from [20, Theorem 1.5] that $\kappa(x,y)=\frac{2+\alpha}{d}$ holds whenever $\alpha=0$ and $\beta\geq 2$. If $\alpha=0$, then the conditions (2) and (3) both lead to $\beta\geq 2$ and hence $\kappa(x,y)=\frac{2}{d}$. Under the condition (1), $\alpha=0$ implies $\beta=1$ and $d>2$. However, by the condition $n<3d-2\alpha$ and Lemma 3.2, we have $d\leq 2\beta=2$, which is a contradiction. That is, under the condition (1), the case $\alpha=0$ is not allowed. In the following proof, we assume that $\alpha>0$.
@@
always holds. For each vertex $v_{i}\in S$, let $x_{i}$ be the number of neighbors of $v_{i}$ in $N_{y}$. By Lemma 3.3, the vertex $v_{i}$ has $d-\alpha-1-x_{i}$ neighbors in $P_{xy}$. If $|P_{xy}|=0$, then we have $x_{i}=d-\alpha-1=|N_{y}|$. Therefore, the inequality (3.3) clearly holds. In the following proof, we assume that $|P_{xy}|>0$, i.e., $n-2d+\alpha>0$.
@@
Assume, for the sake of contradiction, that $|T|\leq b-1$.